OpenStudy (anonymous):

What is the missing number in this pattern? 7, 16, ___, 34, 43

Parth (parthkohli):

Keep adding 9 to get the next term.

OpenStudy (anonymous):

24?

Parth (parthkohli):

7 + 9 = 16 16 + 9 = the missing term the missing term + 9 = 34 34 + 9 = 43

OpenStudy (anonymous):

25

Parth (parthkohli):

What is 16 + 9?

OpenStudy (anonymous):

25

Parth (parthkohli):

Yeah, and that is your missing term.